Question no 3: The average income of the residents of a particular community is your roll no *1000 and sd is your roll no *100:

What is the probability that income of a person, selected at random, is more than average income?

 What is the probability that income of a person, selected at random, is between average income and average income +2000?

What is the probability that income of a person, selected at random, is bbetween average income+1000 and average income +2000?

What is the probability that income of a person, selected at random, is less than average income +1000?


1
Expert's answer
2021-02-02T02:28:22-0500

It is nesessary to asume, that roll no is the number given by a professor or a class instructor to the client. I will provide the common solution, but with no numbers in the answer as it will be impossible. To calculate the answer as a number, please refer to the Standart Normal Table: https://www.ztable.net/


Roll no will be named as "n" later in the solution


Solution:


In this case, we can assume that the distribution is normal with mean equal to "1000n" and standart deviation equal to "100n".


1) We need to calculate the following probability:


"P\{\xi>1000n\}"


Next, we will use the following transformation and Standart Normal Table:


"P\{x_1\le\xi\le x_2\} = P\{(x_1-a)/\sigma \le (\xi-a)/\sigma \le (x_2-a)/\sigma\} ="

"= \Phi((x_2-a)/\sigma) - \Phi((x_1-a)/\sigma)"


"a = 1000n," "\sigma = \sqrt{100n} = 10\sqrt{n}"


Upper boundary is going to be "x_2 = +\infin", "\Phi(+\infin) = 1"(adding or subtracting or divising will yeld no result), then let us calculate the lower boundary, or "x_1 = 1000n":


"\Phi((1000n-1000n)/\sigma) = \Phi(0) = 0.5"


Finally, subtracting, we will get the answer:


"1 - 0.5 = 0.5"


2) We need to find the probability that income of a person is between average and average + 2000. It can be represented as:


"P(E(\xi)\le\xi\le E(\xi) + 2000) = P(1000n\leq \xi \le 1000n+2000)"


Using the transformation from 1):


"P(1000n\leq \xi \le 1000n+2000) = \Phi((1000n + 2000 - 1000n)/(10\sqrt n)) -"

"- \Phi((1000n-1000n)/(10\sqrt n)) = \Phi(2000/(10\sqrt n)) - \Phi(0) ="

"= \Phi(2000/(10\sqrt n)) - 0.5"


3) Following the same procedure as in 2) but for the different interval:


"P\{1000n +1000 \le \xi \le 1000n + 2000\} = \Phi((1000n + 2000 - 1000n)/(10\sqrt n)) -"

"\Phi((1000n + 1000)/10\sqrt n) = \Phi(2000/(10\sqrt n)) - \Phi(1000/ (10\sqrt n))"


4) Now we need to calculate the following probability:


"P(\xi < 1000n + 1000)", for transformation lower boundary will be "- \infin", "\Phi(-\infin) = 0"

So, the result will be:


"\Phi((1000n+1000-1000n)/(10\sqrt n)) = \Phi(1000/(10\sqrt n))"


Answer:

1) 0.5

2) "\Phi(2000/(10\sqrt n)) - 0.5"

3) "\Phi(2000/(10\sqrt n)) - \Phi(1000/ (10\sqrt n))"

4) "\Phi(1000/ (10\sqrt n))"